kurt gets $6.00 a week for doing chores. How many weeks will it take him to save $50.00 if he already has $10.00 saved?
Answer by checkley79(3341)

10+6X=50
6X=50-10
6X=40
X=40/6
X=6.67 WEEKS ANS.